What is dynamic mass of the photon of frequency vAnd si unit of electrical flux

Grade:12

1 Answers

Arun
6 years ago
Dear Gagan
E = h* v
And
E = m *c²
Hence
m c² = h* v
m = h* v/c²
Regards
Arun (askIITians forum expert)